### 09 Mar 2026 v1

# On the stable Hopf invariant

Abstract.
We provide a simplified approach to the the stable Hopf invariant. We provide short elementary proofs of the Cartan Formula, the Composition Formula, and the Transfer formula. In addition, when $\pi$ is a discrete group, we show how to extend these results to the stable category of $\pi$-spaces.
1991 Mathematics Subject Classification:
Primary: 55P91, 55P92, 55S91.
1. Introduction
Heinz Hopf introduced what is now called the Hopf invariant in his 1931 work on maps between spheres as a way to distinguish homotopy classes of maps that are invisible to homology [7]. In the 1950s, this notion was generalized by James [8, pp. 192–193], who defined invariants associated to desuspending maps of based spaces $\Sigma A\to\Sigma B$. For each positive integer $n$, the $n$th James–Hopf invariant is an operation on homotopy classes
$\gamma_{n}\colon[\Sigma A,\Sigma B]\longrightarrow[\Sigma A,\Sigma B^{[n]}]$
which is natural in both $A$ and $B$, where $B^{[n]}$ denotes the $n$-fold smash product of $B$ with itself.
Boardman and Steer [1] instead consider the suspended invariants
$\lambda_{n}=\Sigma^{n-1}\gamma_{n}\colon[\Sigma A,\Sigma B]\longrightarrow[\Sigma^{n}A,\Sigma^{n}B^{[n]}]$
and show that they are axiomatically characterized by certain properties encoded in what they call a Hopf ladder. A Hopf ladder is a collection of such natural transformations satisfying $\lambda_{1}=\mathrm{id}$, $\lambda_{n}(\Sigma f)=0$ for $n\geq 2$, and the Cartan formula
$\lambda_{n}(f+g)=\sum_{i+j=n}\lambda_{i}(f)\cdot\lambda_{j}(g)\,.$
In the 1970s, the Segal-Snaith Hopf invariants $h_{n}$ were introduced using the approximation theorem in combination with the Snaith splitting [14], [10], [2], [3], [13]. Consider the infinite loop space $Q(B)=\Omega^{\infty}\Sigma^{\infty}B$ for a based space $B$. The Snaith splitting provides a decomposition of the suspension spectrum
$\Sigma^{\infty}Q(B)\simeq\bigvee_{n\geq 0}\Sigma^{\infty}D_{n}(B)\,$,
where $D_{n}(B):=B^{[n]}\wedge_{\Sigma_{n}}({E\Sigma_{n}}_{+})$ is the $n$-adic construction, i.e. the reduced homotopy orbits of the symmetric group $\Sigma_{n}$ acting on the $n$-fold smash product $B^{[n]}$. If we compose the Snaith splitting with projection onto the $n$th summand and take the adjoint, we obtain a map
$Q(B)\to Q(D_{n}(B))\,.$
The latter induces the operation
$h_{n}\colon\{A,B\}\to\{A,D_{n}(B)\}\,$
by applying homotopy classes [A,][A,{-}]. In the display, $\{A,B\}=[A,Q(B)]$ is the abelian of homotopy classes of stable maps $A\to B$. As far as I am aware, a Hopf-ladder type axiomatic characterization of the invariants $h_{n}$ is currently unknown.
In this paper we restrict our attention to the case $n=2$. In this case there is a different construction of an operation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
which avoids the Snaith splitting (I first learned of such a construction from Andrew Ranicki). Our main goal is to develop and analyze the fundamental properties of $H$.
Let
$E\colon[A,B]\to\{A,B\}$
be the stabilization map, i.e., the passage from unstable homotopy classes to stable homotopy classes. If $f,g\in\{A,B\}$ then the cup product [1]
$f\cup g\in\{A,B\wedge B\}$
is the stable composition
$A@>\Delta_{A}>>A\wedge A@>f\wedge g>>B\wedge B\,.$
The symmetrized cup product
$f\cup_{2}g\in\{A,D_{2}(B)\}$
is the composition
$A@>f\cup g>>B\wedge B\to D_{2}(B\wedge B)$
where the map $B\wedge B\to D_{2}(B\wedge B)$ is the homotopy quotient by the cyclic group $\mathbb{Z}_{2}=\Sigma_{2}$.
We are now in position to state the main result.
Theorem A.
There is an operation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
called the stable Hopf invariant that is natural in both $A$ and $B$ and which satisfies:
  1. (i).
    (Normalization). $H(E(f))=0$;
  2. (ii).
    (Cartan Formula). $H(f+g)=H(f)+H(g)+f\cup_{2}g$;
  3. (iii).
    (Transfer Formula). $\operatorname{tr}H(f)=f\cup f-\Delta_{B}\circ f$;
  4. (iv).
    (Composition Formula). $H(g\circ f)=H(g)\circ f+D_{2}(g)\circ H(f)$.
Remark 1.1.
I am not aware of any place in the literature where the above results appear in the current form. However, with some effort they can be extracted from the properties of the geometric Hopf invariant of Crabb and Ranicki [4].
There are some key differences between our approach and that of Crabb and Ranicki: They fix a $\mathbb{Z}_{2}$-representation $V$ and define the geometric Hopf invariant as a natural transformation
$h_{V}\colon[S^{V}\wedge A,S^{V}\wedge B]_{\mathbb{Z}_{2}}\to[\Sigma S(\alpha\otimes V)_{+}\wedge S^{V}\wedge A,S^{(\alpha+1)\otimes V}\wedge B]_{\mathbb{Z}_{2}}\,$,
where the domain is the set of homotopy classes of $\mathbb{Z}_{2}$-equivariant maps $S^{V}\wedge A\to S^{V}\wedge B$. In the display, $S^{V}$ denotes the one-point compactification of $V$, $\alpha$ is the sign representation, and $S(\alpha\otimes V)$ is the unit sphere.
By constrast, the source and target of the operation $H$ do not involve a choice of representation and are expressed in terms of unequivariant stable homotopy classes. Even though $H$ is phrased in non-equivariant language, its actual construction requires passing through the $\mathbb{Z}_{2}$-equivariant stable category. The operations $H$ and $h_{V}$ are related by stabilization by representations and the application of a transfer map.
1.1. Relation to the Segal-Snaith operation
Let $h:=h_{2}$ be the second Segal-Snaith Hopf invariant.
Theorem B.
The operations $h,H\colon\{A,B\}\to\{A,D_{2}(B)\}$ coincide.
Suppose that $A$ is a CW complex of dimension $s$ and $Y$ is $N$-connected. Then stable range occurs when $s\leq 2r+1$; in this case $E\colon[A,B]\to\{A,B\}$ is surjective by the Freudenthal suspension theorem. The metastable range occurs when $s\leq 3r+1$. It is well-known that the Segal-Snaith Hopf invariant is the complete obstruction to destabilization in the metastable range [11].
Corollary C.
Assume $s\leq 3r+1$. Let $f\in\{A,B\}$ be a homotopy class. Then $H(f)=0$ if and only if $f=E(f^{\prime})$ for some $f^{\prime}\in[X,Y]$.
1.2. Extension to $\pi$-spaces
For applications to surgery theory, it is useful to generalize the above to the $\pi$-equivariant setting, where $\pi$ is any discrete group. Let $A$ and $B$ be based $\pi$-CW complexes such that the action is free away from the basepoint. In this instance, we have an equivariant stable Hopf invariant
$H^{\pi}\colon\{A,B\}_{\pi}\to\{A,D_{2}(B)\}_{\pi}$
where $\{A,B\}_{\pi}:=[A,Q(B)]_{\pi}$ denotes the abelian group of homotopy classes of stable equivariant maps $A\to B$.
Addendum D.
The equivariant stable Hopf invariant $H^{\pi}$ possesses properties (i)-(iv). Furthemore, the $\pi$-equivariant analog of Corollary C holds: In the metastable range, $f\in\{A,B\}_{\pi}$ destabilizes to $A,B]_{\pi}$ if and only if $H^{\pi}(f)$ is trivial.
Outline.
Section 2 introduces $\mathbb{Z}_{2}$-equivariant stable homotopy from a low-tech point-of-view. Section 3 recalls the tom Dieck splitting in the case of the group $\mathbb{Z}_{2}$. In Section 4 we provide a construction of the stable Hopf invariant. Section 5 concerns the proof of Theorem A and Addendum D. The proof of Theorem B appears in Section 6. Section 7 contains the proof of Addendum D.
2. $\mathbb{Z}_{2}$-equivariant stable homotopy
Let $T$ be the category of compactly generated weak Hausdorff spaces and let $T_{\ast}$ be the corresponding category of based spaces. We let $T_{\ast}(\mathbb{Z}_{2})$ be the category of based spaces with (left) $\mathbb{Z}_{2}$-action. If $X$ is an unbased $\mathbb{Z}_{2}$-space, we let $X_{+}$ denote the corresponding based one obtained from $X$ by taking a disjoint basepoint.
For based $\mathbb{Z}_{2}$-spaces $A,B$, the function space of unequivariant based maps
$\operatorname{map}_{\ast}(A,B)$
is equipped with a preferred $\mathbb{Z}_{2}$-action, where the action is defined by conjugating functions.
Up to isomorphism, there exactly two irreducible real orthogonal representations of $\mathbb{Z}_{2}$ of positive dimension. These are the trivial representation $1$ and the sign representation $\alpha$. We choose a complete universe $\mathscr{U}$ for $\mathbb{Z}_{2}$ given by the direct sum of countably many copies $1$ and $\alpha$. Then every finite dimensional subrepresentation of $\mathscr{U}$ is a direct sum of the form
$s\alpha+t\,$,
for non-negative integers $s$ and $t$.
If $V\subset\mathscr{U}$ is a finite dimensional subrepresentation, then we let $S^{V}$ denote its one-point compactification. The latter is a sphere of dimension $\dim V$ equipped with a based $\mathbb{Z}_{2}$-action.
Suppose that $Y$ is a based $\mathbb{Z}_{2}$-CW complex. We give $S^{V}\wedge Y$ the diagonal action. The based function space
$\Omega^{V}\Sigma^{V}(Y):=\operatorname{map}_{\ast}(S^{V},S^{V}\wedge Y)$
is equipped with a $\mathbb{Z}_{2}$-action. Then the colimit
$Q_{\mathbb{Z}_{2}}(Y):=\operatorname*{colim}_{V\in\mathscr{U}}\Omega^{V}\Sigma^{V}(Y)$
is equipped with a $\mathbb{Z}_{2}$-action. The latter is the zeroth space of a $\mathbb{Z}_{2}$-equivariant spectrum
$\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y)$
whose $V$-th space is $Q_{\mathbb{Z}_{2}}(S^{V}\wedge Y)$.
Definition 2.1.
Let $X$ be a based $\mathbb{Z}_{2}$-CW complex. Set
$\{X,Y\}_{\underline{\mathbb{Z}_{2}}}:=\operatorname*{colim}_{V\subset\mathscr{U}}\pi_{0}(\operatorname{map}_{\ast}(S^{V}\wedge X,S^{V}\wedge Y)^{\mathbb{Z}_{2}})\,.$
Remark 2.2.
Alternatively,
$\{X,Y\}_{\underline{\mathbb{Z}_{2}}}=\pi_{0}(\operatorname{map}_{\ast}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}})$
where $\operatorname{map}_{\ast}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}}$ is the based mapping space of $C$-equivariant maps $X\to Q_{\mathbb{Z}_{2}}(Y)$.
Definition 2.3 ($\mathbb{Z}_{2}$-equivariant Stable Maps).
Suppose that $X$ and $Y$ are based $\mathbb{Z}_{2}$-CW complexes. An equivariant stable map $X\to Y$ is an equivariant map $X\to Q_{\mathbb{Z}_{2}}(Y)$.
Stable maps can be composed and form an $\infty$-category. If $X$ is finite, then there is a finite dimensional representation $V\subset\mathscr{U}$ such that a stable map is represented by an equivariant map $S^{V}\wedge X\to S^{V}\wedge Y$, where two such maps agree if they coincide in the colimit of the mapping space.
3. The tom Dieck splitting for $\mathbb{Z}_{2}$
Let $X$ be free unbased $\mathbb{Z}_{2}$-CW complex. We let $X/\mathbb{Z}_{2}$ be the orbit space. Then one has an Adams isomorphism [9, §II.7]
$\Sigma^{\infty}((X/\mathbb{Z}_{2})_{+})@>\simeq>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}X_{+})^{\mathbb{Z}_{2}}\,$,
where the target is the categorical fixed points of $\mathbb{Z}_{2}$ acting on the genuine $\mathbb{Z}_{2}$-spectrum $\Sigma_{\mathbb{Z}_{2}}(X_{+})$. The Adams isomorphism is induced by the stable transfer map for the regular covering space
$X\to X/\mathbb{Z}_{2}\,.$
More generally, suppose that $Y$ is an arbitrary based $\mathbb{Z}_{2}$-CW complex. If we replace $Y$ by $Y\times E\mathbb{Z}_{2}$, then one has a regular covering space of pairs
$(Y\times E\mathbb{Z}_{2},E\mathbb{Z}_{2})\to(Y\times_{\mathbb{Z}_{2}}E\mathbb{Z}_{2},B\mathbb{Z}_{2})$
and therefore a relative transfer map of pairs which induces on quotients an equivalence
$\Sigma^{\infty}(Y_{h\mathbb{Z}_{2}})@>>>\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y\wedge({E\mathbb{Z}_{2}}_{+}))^{\mathbb{Z}_{2}}\,$,
where $Y_{h\mathbb{Z}_{2}}:=Y\wedge_{\mathbb{Z}_{2}}({E\mathbb{Z}_{2}}_{+})$ is the reduced Borel construction.
Remark 3.1.
The classical transfer map $\operatorname{tr}\colon\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to\Sigma^{\infty}Y$ of the covering space $(Y\times E\mathbb{Z}_{2},E\mathbb{Z}_{2})\to(Y\times_{\mathbb{Z}_{2}}E\mathbb{Z}_{2},B\mathbb{Z}_{2})$ is the composition
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}@>\subset>>\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y)\simeq\Sigma^{\infty}(Y)\,.$
With $Y$ as above, there is always a split cofiber sequence of spectra
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}@>\eta>>\Sigma^{\infty}(Y^{\mathbb{Z}_{2}})\,.$
The map $\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$ is a composition:
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}@>\simeq>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y\wedge({E\mathbb{Z}_{2}}_{+})))^{\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$
where the first map is the Adams isomorphism and the second map is induced by collapsing $E\mathbb{Z}_{2}$ to a point.
The splitting map $\Sigma^{\infty}(Y^{\mathbb{Z}_{2}})\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$ is induced by the inclusion $Y^{\mathbb{Z}_{2}}\to Y$, more precisely, it is adjoint to the map
$Y^{\mathbb{Z}_{2}}\to(Q_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$
which sends a a fixed point of $y\in Y^{\mathbb{Z}_{2}}$ to the point represented by the equivariant map $S^{0}=S^{0}\wedge(\{y\}_{+})\to S^{0}\wedge Y$.
3.1. Specialization
Let $B$ be a based CW complex. We will be interested in applying the above observation to the based $\mathbb{Z}_{2}$-CW complex
$Y=B\wedge B\,$,
where the action of $\mathbb{Z}_{2}$ is given by permuting the factors of the smash product. In this case the split cofiber sequence is
$\Sigma^{\infty}D_{2}(B)@>\iota_{B}>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}\to\Sigma^{\infty}B\,$,
since $(B\wedge B)^{\mathbb{Z}_{2}}=B$. The splitting map for the cofiber sequence is induced by the reduced diagonal $\Delta_{B}\colon B\to B\wedge B$.
The inclusion $\mathbb{Z}_{2}\to E\mathbb{Z}_{2}$ and the constant map $\mathbb{Z}_{2}\to\ast$ induce a commutative square
${\Sigma^{\infty}(B\wedge B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{({-})_{2}}$${(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B\wedge{\mathbb{Z}_{2}}_{+}))^{\mathbb{Z}_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Sigma^{\infty}(D_{2}(B))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{B}}$${(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}}$
The left vertical map $({-})_{2}$ is given by
$B\wedge B=(B\wedge B)\wedge_{\mathbb{Z}_{2}}({\mathbb{Z}_{2}}_{+})@>>>(B\wedge B)\wedge_{\mathbb{Z}_{2}}({E\mathbb{Z}_{2}}_{+})=D_{2}(B)\,.$
The top map of square is an equivalence by the Adams isomorphism. Taking the corresponding infinite loop spaces, the composition
(1) $Q(B\wedge B)@>({-})_{2}>>Q(D_{2}(B))@>\iota_{B}>>(Q_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}$
is defined as follows: Give $S^{n}$ the trivial $\mathbb{Z}_{2}$-action. If $\alpha\colon S^{n}\to S^{n}\wedge(B\wedge B)$ is a map representing a point of $Q(B\wedge B)$, then we form the symmetrization
$(1+\tau)\circ\alpha:=\alpha+\tau\circ\alpha\colon S^{n}\to S^{n}\wedge(B\wedge B)\,$,
where $\tau\colon S^{n}\wedge B\wedge B\to S^{n}\wedge B\wedge B$ is the map switches the $B$ factors. Here addition is defined in the usual way by
$S^{n}@>\text{pinch}>>S^{n}\vee S^{n}@>\alpha\vee\tau\circ\alpha>>S^{n}\wedge(B\wedge B)\vee S^{n}\wedge(B\wedge B)@>\text{fold}>>S^{n}\wedge(B\wedge B)\,.$
Then $(1+\tau)\circ\alpha$ is $\mathbb{Z}_{2}$-equivariant so it defines a point of $(Q_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}$, and we have
(2) $\iota_{B}\circ(\alpha)_{2}=(1+\tau)\circ\alpha\,.$
4. Definition of the stable Hopf invariant
Given a stable map $f\colon A\to B$, i.e., a map $A\to Q(B)$, we may associate two equivariant stable maps $\Delta_{B}\circ f,(f\wedge f)\circ\Delta_{B}\colon A\to B\wedge B$.
The map $\Delta_{B}\circ f$ is given by the composition
$S^{n}\wedge A@>f>>S^{n}\wedge B@>\Delta_{B}>>S^{n}\wedge B\wedge B\,.$
For the second map, take $f\wedge f$ and shuffle To obtain an equivariant map
$S^{n}\wedge S^{n}\wedge A\wedge A\cong S^{n}\wedge A\wedge S^{n}\wedge A@>f\wedge f>>S^{n}\wedge B\wedge S^{n}\wedge B\cong S^{n}\wedge S^{n}\wedge B\wedge B\,.$
where $S^{n}\wedge S^{n}$ is given the action that switches factors. Next, note that $S^{n}\wedge S^{n}\cong S^{n\alpha+n}$, where $\alpha$ is the sign representation, and $n\alpha+n$ is notation for the direct sum of $n$-copies of $\alpha$ with $n$-copies of the trivial representation.
Making this identification and taking the composition
$S^{n\alpha+n}\wedge A@>\Delta_{A}>>S^{n\alpha+n}\wedge A\wedge A@>f\wedge f>>S^{n\alpha+n}\wedge B\wedge B$
yields the desired equivariant stable map $(f\wedge f)\circ\Delta_{B}$.
By taking adjunctions, we have constructed two maps $A\to Q_{\mathbb{Z}_{2}}(B\wedge B)^{\mathbb{Z}_{2}}$. Moreover, it is readily checked that the diagram
$A\,\raisebox{-2.58334pt}{\hbox{\set@color\hskip 46.46922pt\hskip-41.72163pt\hbox{\set@color$@>>\phantom{aa}\Delta_{B}\circ f\phantom{aa}>$}\hskip-41.72163pt\hskip-46.46922pt\raisebox{7.44444pt}{\hbox{\set@color$@>(f\wedge f)\circ\Delta_{B}>>$}}\hskip-46.46922pt\hskip 46.46922pt}}\,Q_{\mathbb{Z}_{2}}(B\wedge B)^{\mathbb{Z}_{2}}@>\eta>>Q(B^{\mathbb{Z}_{2}})$
commutes, i.e., $\eta$ coequalizes the maps $\Delta_{B}\circ f,(f\wedge f)\circ\Delta_{B}$. It follows that the homotopy class of the difference
$(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\in\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}$
is in the image of the split injection
$\iota_{B}\colon\{A,D_{2}(B)\}\to\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}\,.$
Definition 4.1.
The stable Hopf invariant is the natural transformation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
such that $H(f)\in\{A,D_{2}(B)\}$ is the unique element such that
(3) $\iota_{B}H(f)=(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\,.$
5. Proof of Theorem A
5.1. The cup product again
Let $A$ and $B$ be CW complexes. For stable maps $f,g\colon A\to B$, the cup product
(4) $f\cup g\in\{A,B\wedge B\}$
is defined by the composition of the equivariant stable maps $f\wedge g$ and $\Delta_{A}$, i.e.,
$f\cup g:=(f\wedge g)\circ\Delta_{A}\,.$
If $A$ is finite, then $f\cup g$ is given as follows: Let $f\colon\Sigma^{j}A\to\Sigma^{j}B$ and $g\colon\Sigma^{j}A\to\Sigma^{j}B$ be representatives. Then $f\cup g$ is given by the composition
$\Sigma^{2j}A@>\Sigma^{2j}\Delta_{A}>>\Sigma^{2j}(A\wedge A)=\Sigma^{j}A\wedge\Sigma^{j}A@>f\wedge g>>\Sigma^{j}B\wedge\Sigma^{j}B=\Sigma^{2j}(B\wedge B)\,.$
The cup product induces a pairing
$\cup\colon\{A,B\}\times\{A,B\}\to\{A,B\wedge B\}\,.$
The proof of the following result is straightforward and we leave its details to the reader.
Lemma 5.1.
The cup product satisfies the following identities:
  1. (i).
    $g\cup f=\tau\circ(f\cup g)$, where $\tau\colon B\wedge B\to B\wedge B$ is the twist map;
  2. (ii).
    $(f+f^{\prime})\cup g=f\cup g+f^{\prime}\cup g$.
Definition 5.2.
For equivariant stable maps $f,g\colon A\to B$, the symmetrized cup product
$f\cup_{2}g:=(f\cup g)_{2}\in\{A,D_{2}(B)\}$
is the stable composition
$A@>f\cup g>>B\wedge B@>({-})_{2}>>D_{2}(B)\,.$
Proof of Theorem A.
As remarked above, $\iota_{B}\colon\{A,D_{2}(B)\}\to\{A,B\wedge B)\}_{\underline{\mathbb{Z}_{2}}}$ is a split injection. Recall that $H(f)$ is defined using the identity (3). For the Normalization property, it suffices to note that when $f\colon A\to B$ is an unstable map, then the square
[…]
commutes. It follows that the difference $\iota_{B}H(f):=(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\in\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}$ is trivial.
The Transfer Formula is also an easy consequence of the defining identity (3), since the composition
$\{A,D_{2}(B)\}@>\iota_{B}>>\{A,B\wedge B)\}_{\underline{\mathbb{Z}_{2}}}\to\{A,B\wedge B\}$
gives the transfer, where the second displayed map is the forgetful homomorphism.
The Composition Formula also follows from (3): If $f\colon A\to B$ and $g\colon B\to C$, then
$\iota_{C}H(g\circ f)$ $=(g\circ f)\wedge(g\circ f)\circ\Delta_{A}-\Delta_{C}\circ g\circ f\,$,
$=(g\wedge g)\circ(f\wedge f)\circ\Delta_{A}-(\Delta_{C}\circ g)\circ f\,$,
$=(g\wedge g)\circ((f\wedge f)\circ\Delta_{B}-\Delta_{B}\circ f)+(g\wedge g)\circ\Delta_{B}-(\Delta_{C}\circ g))\circ f\,$,
$=(g\wedge g)\circ\iota_{B}H(f)+\iota_{C}H(g)\circ f\,$,
$=\iota_{C}((g\wedge g)\circ H(f)+H(g)\circ f)\,.$
For the Cartan formula, we compute
$\iota_{B}H(f+g)$ $=((f+g)\wedge(f+g))\circ\Delta_{A}-\Delta_{B}\circ(f+g)\,$,
$=(f\wedge f)\circ\Delta_{A}+(f\wedge g)\circ\Delta_{A}+(g\wedge f)\circ\Delta_{A}+(g\wedge g)\circ\Delta_{A})-(\Delta_{B}\circ f+\Delta_{B}\circ g)\,$,
$=((f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f))+((g\wedge g)\circ\Delta_{A}-\Delta_{B}\circ g)+((f\wedge f)\circ\Delta_{A}+(g\wedge g)\circ\Delta_{A})\,$,
$=\iota_{B}H(f)+\iota_{B}H(g)+(f\cup g+g\cup f)\,$,
$=\iota_{B}H(f)+\iota_{B}H(g)+\iota_{B}(f\cup_{2}g)\,$,
$=\iota_{B}(H(f)+H(g)+f\cup_{2}g)\,$,
where we have used the identity $(f\cup g+g\cup f)=(1+\tau)\circ f\cup g=\iota_{B}(f\cup_{2}g)$ arising from (2). ∎
6. Proof of Theorem B
In this paper we provided an alternative construction of a stable Hopf invariant using the tom Dieck splitting. Here we will verify that our construction coincides with the Segal-Snaith Hopf invariant.
Recall from [10] that there is a homotopy functor on based spaces C()C({-}), and a natural transformation
$C(X)@>\eta>>Q(X)\,$,
which is a weak equivalence when $X$ is cofibrant and connected.
The space $C(X)$ is given by a filtration $C_{1}(X)\subset C_{2}(X)\subset\cdots$ and
$\mathop{{\coprod}}_{n\geq 0}(X^{\times n}\times_{\Sigma_{n}}E_{n})/\!\!\sim\,$,
where $E_{n}$ is the space of $n$-disjoint “little cubes” in $\mathbb{R}^{\infty}$. Note that $E_{n}$ is a model for $E\Sigma_{n}$, i.e., a free contractible $\Sigma_{n}$-space, and the $E_{\bullet}$ collectively form the little cubes operad. The equivalence relation is generated by two kinds of operations: The first operation identifies a representative $(x,c)\in X^{\times(n-1)}\times E_{n}$ with $(s_{j}x,c)$, where $s_{j}$ inserts the basepoint in the $j$-factor of $X^{\times n}$ and the second operation identifies $(x,c)$ with $(x,d_{j}(c))$, where $d_{j}(c):E\Sigma_{n}\to E\Sigma_{n-1}$ is given by deleting the $j$-th cube.
The Pontryagin-Thom construction defines an equivariant map
$\theta\colon E_{n}\to Q(\vee_{n}S^{0})\,$
which underlies a morphism of operads $E_{\bullet}\to Q(\vee_{\bullet}S^{0})$. In particular, if $t\in E_{2}$, then $\theta(t)\colon S^{0}\to S^{0}\vee S^{0}$ parametrizes a multiplication on the infinite loop space $Q(X)$: If $f\,g\colon S^{0}\to X$ are stable maps then
$S^{0}@>\theta(t)>>S^{0}\vee S^{0}@>f\vee g>>X\vee X@>\text{fold}>>X$
defines $f+g\in Q(X)$, or more precisely, if we wish to indicate the dependence on $t$, we write $f+_{t}g$.
Furthermore, $Q(X)$ is algebra over $Q(\vee_{\bullet}S^{0})$. Then $\eta\colon C(X)\to Q(X)$ is the composition
$C(X)=C(E_{\bullet};X)\to C(\vee_{\bullet}S^{0};X)\to Q(X)\,$,
where $C(\vee_{\bullet}S^{0};X)$ is defined with $\vee_{\bullet}S^{0}$ in place of $E_{\bullet}$. The map $\eta$ is a weak homotopy equivalence if $X$ is cofibrant and connected. Furthermore, there is a natural equivalence
$\Sigma^{\infty}C(X)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)$
such that the composition
$\Sigma^{\infty}((X^{\times n}\times_{\Sigma_{n}}{E_{n}})_{+})\to\Sigma^{\infty}C(X)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)$
is homotopic to the evident projection.
Proof of Theorem B.
By the above discussion, the operations $h$ and $H$ are induced by natural transformations of spectrum-valued homotopy functors
${\mathbf{h}},{\mathbf{H}}\colon\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)\to\Sigma^{\infty}D_{2}(X)\,.$
Consequently, for every integer $n\geq 1$, we have a pair if natural transformations of homotopy functors
${\mathbf{h}}_{n},{\mathbf{H}}_{n}\colon\Sigma^{\infty}D_{n}(X)\to\Sigma^{\infty}D_{2}(X)\,$
defined respectively by the restrictions of ${\mathbf{h}},{\mathbf{H}}$ to $\Sigma^{\infty}D_{n}(X)$. By construction, the natural transformation $\mathbf{h}_{n}$ is trivial if $n\neq 2$ and $\mathbf{h}_{2}$ is the identity.
Since the functor $X\mapsto\Sigma^{\infty}D_{n}(X)$ is homogeneous of degree $n$, by standard Goodwillie calculus arguments [5], the natural transformations ${\mathbf{H}}_{n}$ are trivial for $n>2$. It therefore suffices to show that ${\mathbf{H}}_{1}$ is trivial and that ${\mathbf{H}}_{2}$ is the identity.
The map $\iota_{X}\circ H\colon Q(X)\to(Q_{\mathbb{Z}_{2}}(X\wedge X))^{\mathbb{Z}_{2}}$ is a difference of two maps. The first of these sends a representative $f\colon S^{j}\to S^{j}\wedge X$ to
$S^{j\alpha+n}=S^{n}\wedge S^{j}@>f\wedge f>>S^{j}\wedge X\wedge S^{j}\wedge X=S^{j\alpha+j}\wedge X\wedge X$
and the second maps $f$ to
$S^{j\alpha+j}=S^{j}\wedge S^{j}@>\text{id}\wedge f>>S^{j}\wedge S^{j}\wedge X@>\text{id}\wedge\Delta_{X}>>S^{j}\wedge S^{j}\wedge X\wedge X=S^{j\alpha+j}\wedge X\wedge X\,.$
Informally, we write $\iota_{X}\circ H(f)=f\wedge f-\Delta_{X}\circ f$.
In particular, if $x\colon S^{0}\to X$ is an unstable based map determined by a choice of point $x\in X$, then $x\wedge x=\Delta_{X}\circ x$. The space $C_{1}(X)=X\times E_{1}$ is identified up to homotopy with $X$. A point of $x\in C_{1}(X)$ amounts to an unstable based map $x\colon S^{0}\to X$. We infer that natural transformation ${\mathbf{H}}_{1}$ is trivial.
It remains to prove that ${\mathbf{H}}_{2}=\text{id}={\mathbf{h}}_{2}$. This will require some preparation. It will be convenient to redefine $D_{2}(X)$ as $X^{[2]}\wedge_{\mathbb{Z}_{2}}({E_{2}}_{+})$. It is well known that the evident map of based spaces $X_{+}\to X$ admits a stable section up to homotopy, i.e., there is a stable map $X\to X_{+}$ such that the stable composite $X\to X_{+}\to X$ is homotopic to the identity. This implies that the induced map $c\colon D_{2}(X_{+})\to D_{2}(X)$ also admits a stable section up to homotopy. Then for any infinite loop space $Z$, the induced homomorphism $c^{\ast}\colon[D_{2}(X),Z]\to[D_{2}(X_{+}),Z]$ is injective. Furthermore,
$D_{2}(X_{+})=(X_{+})^{[2]}\wedge_{\mathbb{Z}_{2}}({E_{2}}_{+})\cong(X^{\times 2}\times_{\mathbb{Z}_{2}}{E_{2}})_{+}\,.$
Consider the composition
(5) $D_{2}(X_{+})@>c>>D_{2}(X)@>\hat{\mathbf{h}}_{2}>>Q(D_{2}(X))@>\iota_{X}>>Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}\,$,
where $\hat{\mathbf{h}}_{2}$ is the inclusion map.
If $((x,y),t)\in X^{\times 2}\times E_{2}$ is a point, we write $(x,y),t]\in X^{\times 2}\times_{\mathbb{Z}_{2}}E_{2}$ for the projection of $((x,y),t)$ to its $\mathbb{Z}_{2}$-orbit. Then (5) is given by
(6) $(x,y),t]\mapsto x\wedge y+y\wedge x\,$,
where again we write $x,y\colon S^{0}\to X$ for the based maps defined by the points $x,y\in X$.
Similarly, using definition of $H$, we see by direct calculation that the composition
(7) $D_{2}(X_{+})@>c>>D_{2}(X)@>\hat{\mathbf{H}}_{2}>>Q(D_{2}(X))@>\iota_{X}>>Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$
(in which $\hat{\mathbf{H}}_{2}$ is adjoint to ${\mathbf{H}}_{2}$) is given by
(8) $(x,y),t]\mapsto(x+y)\wedge(x+y)-((x\wedge x)+y\wedge y))\,.$
The map $D(X_{+})\to Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$ defined by $(x,y),t]\mapsto(x+y)\wedge(x+y)$ is homotopic to the sum of three terms
$(x,y),t]\mapsto x\wedge y+(x\wedge y+y\wedge x)+y\wedge y\,$,
where each of the terms defines a map $D(X_{+})\to Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$ in its own right. Therefore, (8) is homotopic to the signed sum of five maps
$(x,y),t]\mapsto x\wedge x+(x\wedge y+y\wedge x)+y\wedge y-x\wedge x-y\wedge y\,$,\\
and by homotopy commutativity, the latter is homotopic to the map (6). We infer that the (5) and (7) are homotopic.
As $\iota_{X}$ is a homotopy retract it follows that $\hat{\mathbf{H}}_{2}\circ c$ is homotopic to $\hat{\mathbf{h}}_{2}\circ c$, i.e., on homotopy classes $c^{\ast}(\hat{\mathbf{H}}_{2})=c^{\ast}(\hat{\mathbf{h}}_{2})$. Since $c^{\ast}$ is injective, we conclude that ${\mathbf{H}}_{2}$ is homotopic to ${\mathbf{h}}_{2}$. ∎
7. Proof of Addendum D
Let $T$ be the Quillen model category of compactly generated weak Hausdorff spaces [12]. A weak equivalence of $T$ is a weak homotopy equivalence. A fibration is a Serre fibration. A cofibration is a map which possesses the left lifting property with respect to the acyclic fibrations.
Let $T_{\ast}(\pi)$ be the category based left $\pi$-spaces and $\pi$-equivariant maps; this category has a zero object, i.e., the space $\ast$ consisting of a single point. Declare a map to be a weak equivalence (fibration) if and only if it is a weak equivalence (resp. fibration) upon application of the forgetful functor $T_{\ast}(\pi)\to T$. A map is a cofibration if it possesses the left lifiting property with respect to the acyclic fibrations. Then with respect to these choices, $T_{\ast}(\pi)$ is a model category by application of [6, thm. 11.6.1]. Note that the cofibrant objects of $T_{\ast}(\pi)$ are retracts of objects built up from a point by attaching free cells, where a free cell of dimension $j$ is $D^{j}\times\pi$. In particular, a free $\pi$-CW complex (in the base sense) is cofibrant. Note that every object is fibrant.
For cofibrant objects $X,Y\in T_{\ast}(\pi)$, let
$X,Y]_{\pi}=\hom_{\text{ho}T_{\ast}(\pi)}(X,Y)\,.$
Note that $Q_{\mathbb{Z}_{2}}(Y)$ has the structure of an object of $T_{\ast}(\pi\times\mathbb{Z}_{2})$. Suppose in addition $X$ has the structure of a based $(\pi\times\mathbb{Z}_{2})$-CW complex which is $\pi$-free after forgetting the $\mathbb{Z}_{2}$-action. In this case, we set
$\{X,Y\}_{\pi\times\underline{\mathbb{Z}_{2}}}:=\pi_{0}(\operatorname{map}_{T_{\ast}(\pi)}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}})\,$,
where $\operatorname{map}_{T_{\ast}(\pi)}(A,B)$ denotes the mapping space of based equivariant maps $A\to B$.
Proof of Addendum D.
With respect to the above conventions, the equivariant stable Hopf invariant is defined in the same way as when $\pi$ is trivial: For a cofibrant $\pi$-space $Y$, the tom Dieck fiber sequence
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}\to\Sigma^{\infty}Y^{\mathbb{Z}_{2}}$
is a fiber sequence of naive spectra with $\pi$-action which splits equivariantly making use of the inclusion $Y^{C^{2}}\to Y$ If we set $Y=B\wedge B$ where $B\in T_{\ast}(\pi)$ is a cofibrant object, we see then the construction of §4 yields
$H^{\pi}\colon\{A,B\}_{\pi}\to\{A,D_{2}(B)\}_{\pi}\,.$
The proof that properties (i)-(iv) hold is exactly the same as the proof of Theorem A above, where the maps and homotopy classes are to be interpreted equivariantly.
It remains to show that in the metastable range, a homotopy class $f\in\{A,B\}_{\pi}$ destabilizes if and only if $H^{\pi}(f)=0$. We will provide a crude sketch of the argument and leave the details to the reader. The Snaith splitting is natural, [^1] More precisely, the proofs that appear in [3] and [5] are natural. so the weak homotopy equivalence
$\Sigma^{\infty}Q(B)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(B)$
is also $\pi$-equivariant. It follows that a $\pi$-equivariant version $h^{\pi}$ of the second Segal-Snaith Hopf invariant $h$ is defined. Moreover, the sequence
$B\to Q(B)@>h^{\pi}>>Q(D_{2}(B))$
is an equivariant fiber sequence in the metastable range by the Blakers-Massey excision theorem, where the map $h^{\pi}$ is induced by projection onto the second summand of the Snaith splitting. Then the equivariant stable Hopf invariant $h^{\pi}$ is the the effect of applying $A,{-}]_{\pi}$ to the map of the same name. We infer that $h^{\pi}$ yields the total obstruction to destabilization in the metastable range. Lastly, the proof we gave that $H=h$ is valid in the equivariant setting, and we conclude that $H^{\pi}=h^{\pi}$. ∎