### 10 Mar 2026 v2

# On the stable Hopf invariant

Abstract.
We provide a simplified approach to the stable Hopf invariant. We provide short elementary proofs of the Cartan Formula, the Composition Formula, and the Transfer formula. In addition, when $\pi$ is a discrete group, we show how to extend these results to the stable category of $\pi$-spaces.
1991 Mathematics Subject Classification:
Primary: 55P91, 55P92, 55S91.
1. Introduction
Heinz Hopf introduced what is now called the Hopf invariant in his 1931 work on maps between spheres as a way to distinguish homotopy classes of maps that are invisible to homology [8]. In the 1950s, this notion was generalized by James [9, pp. 192–193], who defined invariants associated to desuspending maps of based spaces $\Sigma A\to\Sigma B$. For each positive integer $n$, the $n$th James–Hopf invariant is an operation on homotopy classes
$\gamma_{n}\colon[\Sigma A,\Sigma B]\longrightarrow[\Sigma A,\Sigma B^{[n]}]$
which is natural in both $A$ and $B$, where $B^{[n]}$ denotes the $n$-fold smash product of $B$ with itself.
Boardman and Steer [1] instead consider the suspended invariants
$\lambda_{n}=\Sigma^{n-1}\gamma_{n}\colon[\Sigma A,\Sigma B]\longrightarrow[\Sigma^{n}A,\Sigma^{n}B^{[n]}]$
and show that they are axiomatically characterized by certain properties encoded in what they call a Hopf ladder. A Hopf ladder is a collection of such natural transformations satisfying $\lambda_{1}=\mathrm{id}$, $\lambda_{n}(\Sigma f)=0$ for $n\geq 2$, and the Cartan formula
$\lambda_{n}(f+g)=\sum_{i+j=n}\lambda_{i}(f)\cdot\lambda_{j}(g)\,.$
In the 1970s, the Segal-Snaith Hopf invariants $h_{n}$ were introduced using the approximation theorem in combination with the Snaith splitting [17], [13], [3], [4], [16]. Consider the infinite loop space $Q(B)=\Omega^{\infty}\Sigma^{\infty}B$ for a based space $B$. The Snaith splitting provides a decomposition of the suspension spectrum
$\Sigma^{\infty}Q(B)\simeq\bigvee_{n\geq 0}\Sigma^{\infty}D_{n}(B)\,$,
where $D_{n}(B):=B^{[n]}\wedge_{\Sigma_{n}}({E\Sigma_{n}}_{+})$ is the $n$-adic construction, i.e. the reduced homotopy orbits of the symmetric group $\Sigma_{n}$ acting on the $n$-fold smash product $B^{[n]}$. If we compose the Snaith splitting with projection onto the $n$th summand and take the adjoint, we obtain a map
$Q(B)\to Q(D_{n}(B))\,.$
The latter induces the operation
$h_{n}\colon\{A,B\}\to\{A,D_{n}(B)\}\,$
by applying homotopy classes [A,][A,{-}]. In the display, $\{A,B\}=[A,Q(B)]$ is the abelian of homotopy classes of stable maps $A\to B$. As far as I am aware, a Hopf-ladder type axiomatic characterization of the invariants $h_{n}$ is currently unknown.
In this paper we restrict our attention to the case $n=2$. In this case there is a different construction of an operation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
which avoids the Snaith splitting (I first learned of such a construction from Andrew Ranicki). Our main goal is to develop and analyze the fundamental properties of $H$.
Let
$E\colon[A,B]\to\{A,B\}$
be the stabilization map, i.e., the passage from unstable homotopy classes to stable homotopy classes. If $f,g\in\{A,B\}$ then the cup product [1]
$f\cup g\in\{A,B\wedge B\}$
is the stable composition
$A@>\Delta_{A}>>A\wedge A@>f\wedge g>>B\wedge B\,.$
The symmetrized cup product
$f\cup_{2}g\in\{A,D_{2}(B)\}$
is the composition
$A@>f\cup g>>B\wedge B\to D_{2}(B\wedge B)$
where the map $B\wedge B\to D_{2}(B\wedge B)$ is the homotopy quotient by the cyclic group $\mathbb{Z}_{2}=\Sigma_{2}$.
We are now in position to state the main result.
Theorem A.
There is an operation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
called the stable Hopf invariant that is natural in both $A$ and $B$ and which satisfies:
  1. (i).
    (Normalization). $H(E(f))=0$;
  2. (ii).
    (Cartan Formula). $H(f+g)=H(f)+H(g)+f\cup_{2}g$;
  3. (iii).
    (Transfer Formula). $\operatorname{tr}H(f)=f\cup f-\Delta_{B}\circ f$;
  4. (iv).
    (Composition Formula). $H(g\circ f)=H(g)\circ f+D_{2}(g)\circ H(f)$.
Remark 1.1.
For the Segal-Snaith Hopf invariant, the Cartan formula appears in [2] (for a geometric description, see [10, thm.2.2]; see also [11] for additional properties of the Segal-Snaith Hopf invariant). For the geometric Hopf invariant of Crabb and Ranicki, a version of the Transfer and Composition Formulas appear in [5, prop. 5.33].
Our approach most closely resembles the the one taken by Crabb and Ranicki with some key differences: They fix a $\mathbb{Z}_{2}$-representation $V$ and define the geometric Hopf invariant as a natural transformation
$h_{V}\colon[S^{V}\wedge A,S^{V}\wedge B]_{\mathbb{Z}_{2}}\to[\Sigma S(\alpha\otimes V)_{+}\wedge S^{V}\wedge A,S^{(\alpha+1)\otimes V}\wedge B]_{\mathbb{Z}_{2}}\,$,
where the domain is the set of homotopy classes of $\mathbb{Z}_{2}$-equivariant maps $S^{V}\wedge A\to S^{V}\wedge B$. In the display, $S^{V}$ denotes the one-point compactification of $V$, $\alpha$ is the sign representation, and $S(\alpha\otimes V)$ is the unit sphere.
By constrast, the source and target of the operation $H$ do not involve a choice of representation and are expressed in terms of unequivariant stable homotopy classes. Even though $H$ is phrased in unequivariant language, its actual construction requires passing through the $\mathbb{Z}_{2}$-equivariant stable category. The operations $H$ and $h_{V}$ are related by stabilization by representations and the application of a transfer map.
1.1. Relation to the Segal-Snaith operation
Let $h:=h_{2}$ be the second Segal-Snaith Hopf invariant.
Theorem B.
The operations $h,H\colon\{A,B\}\to\{A,D_{2}(B)\}$ coincide.
Suppose that $A$ is a CW complex of dimension $s$ and $Y$ is $N$-connected. Then stable range occurs when $s\leq 2r+1$; in this case $E\colon[A,B]\to\{A,B\}$ is surjective by the Freudenthal suspension theorem. The metastable range occurs when $s\leq 3r+1$. It is well-known that the Segal-Snaith Hopf invariant is the complete obstruction to destabilization in the metastable range [14].
Corollary C.
Assume $s\leq 3r+1$. Let $f\in\{A,B\}$ be a homotopy class. Then $H(f)=0$ if and only if $f=E(f^{\prime})$ for some $f^{\prime}\in[X,Y]$.
1.2. Extension to $\pi$-spaces
For applications to surgery theory, it is useful to generalize the above to the $\pi$-equivariant setting, where $\pi$ is any discrete group. Let $A$ and $B$ be based $\pi$-CW complexes such that the action is free away from the basepoint. In this instance, we have an equivariant stable Hopf invariant
$H^{\pi}\colon\{A,B\}_{\pi}\to\{A,D_{2}(B)\}_{\pi}$
where $\{A,B\}_{\pi}:=[A,Q(B)]_{\pi}$ denotes the abelian group of homotopy classes of stable equivariant maps $A\to B$.
Addendum D.
The equivariant stable Hopf invariant $H^{\pi}$ possesses properties (i)-(iv). Furthemore, the $\pi$-equivariant analog of Corollary C holds: In the metastable range, $f\in\{A,B\}_{\pi}$ destabilizes to $A,B]_{\pi}$ if and only if $H^{\pi}(f)$ is trivial.
Outline.
Section 2 introduces $\mathbb{Z}_{2}$-equivariant stable homotopy from a low-tech point-of-view. Section 3 recalls the tom Dieck splitting in the case of the group $\mathbb{Z}_{2}$. In Section 4 we provide a construction of the stable Hopf invariant. Section 5 concerns the proof of Theorem A and Addendum D. The proof of Theorem B appears in Section 6. Section 7 contains the proof of Addendum D.
2. $\mathbb{Z}_{2}$-equivariant stable homotopy
Let $T$ be the category of compactly generated weak Hausdorff spaces and let $T_{\ast}$ be the corresponding category of based spaces. We let $T_{\ast}(\mathbb{Z}_{2})$ be the category of based spaces with (left) $\mathbb{Z}_{2}$-action. If $X$ is an unbased $\mathbb{Z}_{2}$-space, we let $X_{+}$ denote the corresponding based one obtained from $X$ by taking a disjoint basepoint.
For based $\mathbb{Z}_{2}$-spaces $A,B$, the function space of unequivariant based maps
$\operatorname{map}_{\ast}(A,B)$
is equipped with a preferred $\mathbb{Z}_{2}$-action, where the action is defined by conjugating functions.
Up to isomorphism, there exactly two irreducible real orthogonal representations of $\mathbb{Z}_{2}$ of positive dimension. These are the trivial representation $1$ and the sign representation $\alpha$. We choose a complete universe $\mathscr{U}$ for $\mathbb{Z}_{2}$ given by the direct sum of countably many copies $1$ and $\alpha$. Then every finite dimensional subrepresentation of $\mathscr{U}$ is a direct sum of the form
$s\alpha+t\,$,
for non-negative integers $s$ and $t$.
If $V\subset\mathscr{U}$ is a finite dimensional subrepresentation, then we let $S^{V}$ denote its one-point compactification. The latter is a sphere of dimension $\dim V$ equipped with a based $\mathbb{Z}_{2}$-action.
Suppose that $Y$ is a based $\mathbb{Z}_{2}$-CW complex. We give $S^{V}\wedge Y$ the diagonal action. The based function space
$\Omega^{V}\Sigma^{V}(Y):=\operatorname{map}_{\ast}(S^{V},S^{V}\wedge Y)$
is equipped with a $\mathbb{Z}_{2}$-action. Then the colimit
$Q_{\mathbb{Z}_{2}}(Y):=\operatorname*{colim}_{V\in\mathscr{U}}\Omega^{V}\Sigma^{V}(Y)$
is equipped with a $\mathbb{Z}_{2}$-action. The latter is the zeroth space of a $\mathbb{Z}_{2}$-equivariant spectrum
$\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y)$
whose $V$-th space is $Q_{\mathbb{Z}_{2}}(S^{V}\wedge Y)$.
Definition 2.1.
Let $X$ be a based $\mathbb{Z}_{2}$-CW complex. Set
$\{X,Y\}_{\underline{\mathbb{Z}_{2}}}:=\operatorname*{colim}_{V\subset\mathscr{U}}\pi_{0}(\operatorname{map}_{\ast}(S^{V}\wedge X,S^{V}\wedge Y)^{\mathbb{Z}_{2}})\,.$
Remark 2.2.
Alternatively,
$\{X,Y\}_{\underline{\mathbb{Z}_{2}}}=\pi_{0}(\operatorname{map}_{\ast}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}})$
where $\operatorname{map}_{\ast}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}}$ is the based mapping space of $C$-equivariant maps $X\to Q_{\mathbb{Z}_{2}}(Y)$.
Definition 2.3 ($\mathbb{Z}_{2}$-equivariant Stable Maps).
Suppose that $X$ and $Y$ are based $\mathbb{Z}_{2}$-CW complexes. An equivariant stable map $X\to Y$ is an equivariant map $X\to Q_{\mathbb{Z}_{2}}(Y)$.
Stable maps can be composed and form an $\infty$-category. If $X$ is finite, then there is a finite dimensional representation $V\subset\mathscr{U}$ such that a stable map is represented by an equivariant map $S^{V}\wedge X\to S^{V}\wedge Y$, where two such maps agree if they coincide in the colimit of the mapping space.
3. The tom Dieck splitting for $\mathbb{Z}_{2}$
Let $X$ be free unbased $\mathbb{Z}_{2}$-CW complex. We let $X/\mathbb{Z}_{2}$ be the orbit space. Then one has an Adams isomorphism [12, §II.7]
$\Sigma^{\infty}((X/\mathbb{Z}_{2})_{+})@>\simeq>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}X_{+})^{\mathbb{Z}_{2}}\,$,
where the target is the categorical fixed points of $\mathbb{Z}_{2}$ acting on the genuine $\mathbb{Z}_{2}$-spectrum $\Sigma_{\mathbb{Z}_{2}}(X_{+})$. The Adams isomorphism is induced by the stable transfer map for the regular covering space
$X\to X/\mathbb{Z}_{2}\,.$
More generally, suppose that $Y$ is an arbitrary based $\mathbb{Z}_{2}$-CW complex. If we replace $Y$ by $Y\times E\mathbb{Z}_{2}$, then one has a regular covering space of pairs
$(Y\times E\mathbb{Z}_{2},E\mathbb{Z}_{2})\to(Y\times_{\mathbb{Z}_{2}}E\mathbb{Z}_{2},B\mathbb{Z}_{2})$
and therefore a relative transfer map of pairs which induces on quotients an equivalence
$\Sigma^{\infty}(Y_{h\mathbb{Z}_{2}})@>>>\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y\wedge({E\mathbb{Z}_{2}}_{+}))^{\mathbb{Z}_{2}}\,$,
where $Y_{h\mathbb{Z}_{2}}:=Y\wedge_{\mathbb{Z}_{2}}({E\mathbb{Z}_{2}}_{+})$ is the reduced Borel construction.
Remark 3.1.
The classical transfer map $\operatorname{tr}\colon\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to\Sigma^{\infty}Y$ of the covering space $(Y\times E\mathbb{Z}_{2},E\mathbb{Z}_{2})\to(Y\times_{\mathbb{Z}_{2}}E\mathbb{Z}_{2},B\mathbb{Z}_{2})$ is the composition
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}@>\subset>>\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y)\simeq\Sigma^{\infty}(Y)\,.$
With $Y$ as above, there is always a split cofiber sequence of spectra
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}@>\eta>>\Sigma^{\infty}(Y^{\mathbb{Z}_{2}})\,.$
The map $\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$ is a composition:
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}@>\simeq>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}(Y\wedge({E\mathbb{Z}_{2}}_{+})))^{\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$
where the first map is the Adams isomorphism and the second map is induced by collapsing $E\mathbb{Z}_{2}$ to a point.
The splitting map $\Sigma^{\infty}(Y^{\mathbb{Z}_{2}})\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$ is induced by the inclusion $Y^{\mathbb{Z}_{2}}\to Y$, more precisely, it is adjoint to the map
$Y^{\mathbb{Z}_{2}}\to(Q_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}$
which sends a a fixed point of $y\in Y^{\mathbb{Z}_{2}}$ to the point represented by the equivariant map $S^{0}=S^{0}\wedge(\{y\}_{+})\to S^{0}\wedge Y$.
3.1. Specialization
Let $B$ be a based CW complex. We will be interested in applying the above observation to the based $\mathbb{Z}_{2}$-CW complex
$Y=B\wedge B\,$,
where the action of $\mathbb{Z}_{2}$ is given by permuting the factors of the smash product. In this case the split cofiber sequence is
$\Sigma^{\infty}D_{2}(B)@>\iota_{B}>>(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}\to\Sigma^{\infty}B\,$,
since $(B\wedge B)^{\mathbb{Z}_{2}}=B$. The splitting map for the cofiber sequence is induced by the reduced diagonal $\Delta_{B}\colon B\to B\wedge B$.
The inclusion $\mathbb{Z}_{2}\to E\mathbb{Z}_{2}$ and the constant map $\mathbb{Z}_{2}\to\ast$ induce a commutative square
${\Sigma^{\infty}(B\wedge B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$$\scriptstyle{({-})_{2}}$${(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B\wedge{\mathbb{Z}_{2}}_{+}))^{\mathbb{Z}_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Sigma^{\infty}(D_{2}(B))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota_{B}}$${(\Sigma^{\infty}_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}}$
The left vertical map $({-})_{2}$ is given by
$B\wedge B=(B\wedge B)\wedge_{\mathbb{Z}_{2}}({\mathbb{Z}_{2}}_{+})@>>>(B\wedge B)\wedge_{\mathbb{Z}_{2}}({E\mathbb{Z}_{2}}_{+})=D_{2}(B)\,.$
The top map of square is an equivalence by the Adams isomorphism. Taking the corresponding infinite loop spaces, the composition
(1) $Q(B\wedge B)@>({-})_{2}>>Q(D_{2}(B))@>\iota_{B}>>(Q_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}$
is defined as follows: Give $S^{n}$ the trivial $\mathbb{Z}_{2}$-action. If $\alpha\colon S^{n}\to S^{n}\wedge(B\wedge B)$ is a map representing a point of $Q(B\wedge B)$, then we form the symmetrization
$(1+\tau)\circ\alpha:=\alpha+\tau\circ\alpha\colon S^{n}\to S^{n}\wedge(B\wedge B)\,$,
where $\tau\colon S^{n}\wedge B\wedge B\to S^{n}\wedge B\wedge B$ is the map switches the $B$ factors. Here addition is defined in the usual way by
$S^{n}@>\text{pinch}>>S^{n}\vee S^{n}@>\alpha\vee\tau\circ\alpha>>S^{n}\wedge(B\wedge B)\vee S^{n}\wedge(B\wedge B)@>\text{fold}>>S^{n}\wedge(B\wedge B)\,.$
Then $(1+\tau)\circ\alpha$ is $\mathbb{Z}_{2}$-equivariant so it defines a point of $(Q_{\mathbb{Z}_{2}}(B\wedge B))^{\mathbb{Z}_{2}}$, and we have
(2) $\iota_{B}\circ(\alpha)_{2}=(1+\tau)\circ\alpha\,.$
4. Definition of the stable Hopf invariant
Given a stable map $f\colon A\to B$, i.e., a map $A\to Q(B)$, we may associate two equivariant stable maps $\Delta_{B}\circ f,(f\wedge f)\circ\Delta_{B}\colon A\to B\wedge B$.
The map $\Delta_{B}\circ f$ is given by the composition
$S^{n}\wedge A@>f>>S^{n}\wedge B@>\Delta_{B}>>S^{n}\wedge B\wedge B\,.$
For the second map, take $f\wedge f$ and shuffle To obtain an equivariant map
$S^{n}\wedge S^{n}\wedge A\wedge A\cong S^{n}\wedge A\wedge S^{n}\wedge A@>f\wedge f>>S^{n}\wedge B\wedge S^{n}\wedge B\cong S^{n}\wedge S^{n}\wedge B\wedge B\,.$
where $S^{n}\wedge S^{n}$ is given the action that switches factors. Next, note that $S^{n}\wedge S^{n}\cong S^{n\alpha+n}$, where $\alpha$ is the sign representation, and $n\alpha+n$ is notation for the direct sum of $n$-copies of $\alpha$ with $n$-copies of the trivial representation.
Making this identification and taking the composition
$S^{n\alpha+n}\wedge A@>\Delta_{A}>>S^{n\alpha+n}\wedge A\wedge A@>f\wedge f>>S^{n\alpha+n}\wedge B\wedge B$
yields the desired equivariant stable map $(f\wedge f)\circ\Delta_{B}$.
By taking adjunctions, we have constructed two maps $A\to Q_{\mathbb{Z}_{2}}(B\wedge B)^{\mathbb{Z}_{2}}$. Moreover, it is readily checked that the diagram
$A\,\raisebox{-2.58334pt}{\hbox{\set@color\hskip 46.46922pt\hskip-41.72163pt\hbox{\set@color$@>>\phantom{aa}\Delta_{B}\circ f\phantom{aa}>$}\hskip-41.72163pt\hskip-46.46922pt\raisebox{7.44444pt}{\hbox{\set@color$@>(f\wedge f)\circ\Delta_{B}>>$}}\hskip-46.46922pt\hskip 46.46922pt}}\,Q_{\mathbb{Z}_{2}}(B\wedge B)^{\mathbb{Z}_{2}}@>\eta>>Q(B^{\mathbb{Z}_{2}})$
commutes, i.e., $\eta$ coequalizes the maps $\Delta_{B}\circ f,(f\wedge f)\circ\Delta_{B}$. It follows that the homotopy class of the difference
$(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\in\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}$
is in the image of the split injection
$\iota_{B}\colon\{A,D_{2}(B)\}\to\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}\,.$
Definition 4.1.
The stable Hopf invariant is the natural transformation
$H\colon\{A,B\}\to\{A,D_{2}(B)\}$
such that $H(f)\in\{A,D_{2}(B)\}$ is the unique element such that
(3) $\iota_{B}H(f)=(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\,.$
5. Proof of Theorem A
5.1. The cup product again
Let $A$ and $B$ be CW complexes. For stable maps $f,g\colon A\to B$, the cup product
(4) $f\cup g\in\{A,B\wedge B\}$
is defined by the composition of the equivariant stable maps $f\wedge g$ and $\Delta_{A}$, i.e.,
$f\cup g:=(f\wedge g)\circ\Delta_{A}\,.$
If $A$ is finite, then $f\cup g$ is given as follows: Let $f\colon\Sigma^{j}A\to\Sigma^{j}B$ and $g\colon\Sigma^{j}A\to\Sigma^{j}B$ be representatives. Then $f\cup g$ is given by the composition
$\Sigma^{2j}A@>\Sigma^{2j}\Delta_{A}>>\Sigma^{2j}(A\wedge A)=\Sigma^{j}A\wedge\Sigma^{j}A@>f\wedge g>>\Sigma^{j}B\wedge\Sigma^{j}B=\Sigma^{2j}(B\wedge B)\,.$
The cup product induces a pairing
$\cup\colon\{A,B\}\times\{A,B\}\to\{A,B\wedge B\}\,.$
The proof of the following result is straightforward and we leave its details to the reader.
Lemma 5.1.
The cup product satisfies the following identities:
  1. (i).
    $g\cup f=\tau\circ(f\cup g)$, where $\tau\colon B\wedge B\to B\wedge B$ is the twist map;
  2. (ii).
    $(f+f^{\prime})\cup g=f\cup g+f^{\prime}\cup g$.
Definition 5.2.
For equivariant stable maps $f,g\colon A\to B$, the symmetrized cup product
$f\cup_{2}g:=(f\cup g)_{2}\in\{A,D_{2}(B)\}$
is the stable composition
$A@>f\cup g>>B\wedge B@>({-})_{2}>>D_{2}(B)\,.$
Proof of Theorem A.
As remarked above, $\iota_{B}\colon\{A,D_{2}(B)\}\to\{A,B\wedge B)\}_{\underline{\mathbb{Z}_{2}}}$ is a split injection. Recall that $H(f)$ is defined using the identity (3). For the Normalization property, it suffices to note that when $f\colon A\to B$ is an unstable map, then the square
[…]
commutes. It follows that the difference $\iota_{B}H(f):=(f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f\in\{A,B\wedge B\}_{\underline{\mathbb{Z}_{2}}}$ is trivial.
The Transfer Formula is also an easy consequence of the defining identity (3), since the composition
$\{A,D_{2}(B)\}@>\iota_{B}>>\{A,B\wedge B)\}_{\underline{\mathbb{Z}_{2}}}\to\{A,B\wedge B\}$
gives the transfer, where the second displayed map is the forgetful homomorphism.
The Composition Formula also follows from (3): If $f\colon A\to B$ and $g\colon B\to C$, then
$\iota_{C}H(g\circ f)$ $=(g\circ f)\wedge(g\circ f)\circ\Delta_{A}-\Delta_{C}\circ g\circ f\,$,
$=(g\wedge g)\circ(f\wedge f)\circ\Delta_{A}-(\Delta_{C}\circ g)\circ f\,$,
$=(g\wedge g)\circ((f\wedge f)\circ\Delta_{B}-\Delta_{B}\circ f)+(g\wedge g)\circ\Delta_{B}-(\Delta_{C}\circ g))\circ f\,$,
$=(g\wedge g)\circ\iota_{B}H(f)+\iota_{C}H(g)\circ f\,$,
$=\iota_{C}((g\wedge g)\circ H(f)+H(g)\circ f)\,.$
For the Cartan formula, we compute
$\iota_{B}H(f+g)$ $=((f+g)\wedge(f+g))\circ\Delta_{A}-\Delta_{B}\circ(f+g)\,$,
$=(f\wedge f)\circ\Delta_{A}+(f\wedge g)\circ\Delta_{A}+(g\wedge f)\circ\Delta_{A}+(g\wedge g)\circ\Delta_{A})-(\Delta_{B}\circ f+\Delta_{B}\circ g)\,$,
$=((f\wedge f)\circ\Delta_{A}-\Delta_{B}\circ f))+((g\wedge g)\circ\Delta_{A}-\Delta_{B}\circ g)+((f\wedge f)\circ\Delta_{A}+(g\wedge g)\circ\Delta_{A})\,$,
$=\iota_{B}H(f)+\iota_{B}H(g)+(f\cup g+g\cup f)\,$,
$=\iota_{B}H(f)+\iota_{B}H(g)+\iota_{B}(f\cup_{2}g)\,$,
$=\iota_{B}(H(f)+H(g)+f\cup_{2}g)\,$,
where we have used the identity $(f\cup g+g\cup f)=(1+\tau)\circ f\cup g=\iota_{B}(f\cup_{2}g)$ arising from (2). ∎
6. Proof of Theorem B
In this paper we provided an alternative construction of a stable Hopf invariant using the tom Dieck splitting. Here we will verify that our construction coincides with the Segal-Snaith Hopf invariant.
Recall from [13] that there is a homotopy functor on based spaces C()C({-}), and a natural transformation
$C(X)@>\eta>>Q(X)\,$,
which is a weak equivalence when $X$ is cofibrant and connected.
The space $C(X)$ is given by a filtration $C_{1}(X)\subset C_{2}(X)\subset\cdots$ and
$\mathop{{\coprod}}_{n\geq 0}(X^{\times n}\times_{\Sigma_{n}}E_{n})/\!\!\sim\,$,
where $E_{n}$ is the space of $n$-disjoint “little cubes” in $\mathbb{R}^{\infty}$. Note that $E_{n}$ is a model for $E\Sigma_{n}$, i.e., a free contractible $\Sigma_{n}$-space, and the $E_{\bullet}$ collectively form the little cubes operad. The equivalence relation is generated by two kinds of operations: The first operation identifies a representative $(x,c)\in X^{\times(n-1)}\times E_{n}$ with $(s_{j}x,c)$, where $s_{j}$ inserts the basepoint in the $j$-factor of $X^{\times n}$ and the second operation identifies $(x,c)$ with $(x,d_{j}(c))$, where $d_{j}(c):E\Sigma_{n}\to E\Sigma_{n-1}$ is given by deleting the $j$-th cube.
The Pontryagin-Thom construction defines an equivariant map
$\theta\colon E_{n}\to Q(\vee_{n}S^{0})\,$
which underlies a morphism of operads $E_{\bullet}\to Q(\vee_{\bullet}S^{0})$. In particular, if $t\in E_{2}$, then $\theta(t)\colon S^{0}\to S^{0}\vee S^{0}$ parametrizes a multiplication on the infinite loop space $Q(X)$: If $f\,g\colon S^{0}\to X$ are stable maps then
$S^{0}@>\theta(t)>>S^{0}\vee S^{0}@>f\vee g>>X\vee X@>\text{fold}>>X$
defines $f+g\in Q(X)$, or more precisely, if we wish to indicate the dependence on $t$, we write $f+_{t}g$.
Furthermore, $Q(X)$ is algebra over $Q(\vee_{\bullet}S^{0})$. Then $\eta\colon C(X)\to Q(X)$ is the composition
$C(X)=C(E_{\bullet};X)\to C(\vee_{\bullet}S^{0};X)\to Q(X)\,$,
where $C(\vee_{\bullet}S^{0};X)$ is defined with $\vee_{\bullet}S^{0}$ in place of $E_{\bullet}$. The map $\eta$ is a weak homotopy equivalence if $X$ is cofibrant and connected. Furthermore, there is a natural equivalence
$\Sigma^{\infty}C(X)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)$
such that the composition
$\Sigma^{\infty}((X^{\times n}\times_{\Sigma_{n}}{E_{n}})_{+})\to\Sigma^{\infty}C(X)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)$
is homotopic to the evident projection.
Proof of Theorem B.
By the above discussion, the operations $h$ and $H$ are induced by natural transformations of spectrum-valued homotopy functors
${\mathbf{h}},{\mathbf{H}}\colon\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(X)\to\Sigma^{\infty}D_{2}(X)\,.$
Consequently, for every integer $n\geq 1$, we have a pair if natural transformations of homotopy functors
${\mathbf{h}}_{n},{\mathbf{H}}_{n}\colon\Sigma^{\infty}D_{n}(X)\to\Sigma^{\infty}D_{2}(X)\,$
defined respectively by the restrictions of ${\mathbf{h}},{\mathbf{H}}$ to $\Sigma^{\infty}D_{n}(X)$. By construction, the natural transformation $\mathbf{h}_{n}$ is trivial if $n\neq 2$ and $\mathbf{h}_{2}$ is the identity.
Since the functor $X\mapsto\Sigma^{\infty}D_{n}(X)$ is homogeneous of degree $n$, by standard Goodwillie calculus arguments [6], the natural transformations ${\mathbf{H}}_{n}$ are trivial for $n>2$. It therefore suffices to show that ${\mathbf{H}}_{1}$ is trivial and that ${\mathbf{H}}_{2}$ is the identity.
The map $\iota_{X}\circ H\colon Q(X)\to(Q_{\mathbb{Z}_{2}}(X\wedge X))^{\mathbb{Z}_{2}}$ is a difference of two maps. The first of these sends a representative $f\colon S^{j}\to S^{j}\wedge X$ to
$S^{j\alpha+n}=S^{n}\wedge S^{j}@>f\wedge f>>S^{j}\wedge X\wedge S^{j}\wedge X=S^{j\alpha+j}\wedge X\wedge X$
and the second maps $f$ to
$S^{j\alpha+j}=S^{j}\wedge S^{j}@>\text{id}\wedge f>>S^{j}\wedge S^{j}\wedge X@>\text{id}\wedge\Delta_{X}>>S^{j}\wedge S^{j}\wedge X\wedge X=S^{j\alpha+j}\wedge X\wedge X\,.$
Informally, we write $\iota_{X}\circ H(f)=f\wedge f-\Delta_{X}\circ f$.
In particular, if $x\colon S^{0}\to X$ is an unstable based map determined by a choice of point $x\in X$, then $x\wedge x=\Delta_{X}\circ x$. The space $C_{1}(X)=X\times E_{1}$ is identified up to homotopy with $X$. A point of $x\in C_{1}(X)$ amounts to an unstable based map $x\colon S^{0}\to X$. We infer that natural transformation ${\mathbf{H}}_{1}$ is trivial.
It remains to prove that ${\mathbf{H}}_{2}=\text{id}={\mathbf{h}}_{2}$. This will require some preparation. It will be convenient to redefine $D_{2}(X)$ as $X^{[2]}\wedge_{\mathbb{Z}_{2}}({E_{2}}_{+})$. It is well known that the evident map of based spaces $X_{+}\to X$ admits a stable section up to homotopy, i.e., there is a stable map $X\to X_{+}$ such that the stable composite $X\to X_{+}\to X$ is homotopic to the identity. This implies that the induced map $c\colon D_{2}(X_{+})\to D_{2}(X)$ also admits a stable section up to homotopy. Then for any infinite loop space $Z$, the induced homomorphism $c^{\ast}\colon[D_{2}(X),Z]\to[D_{2}(X_{+}),Z]$ is injective. Furthermore,
$D_{2}(X_{+})=(X_{+})^{[2]}\wedge_{\mathbb{Z}_{2}}({E_{2}}_{+})\cong(X^{\times 2}\times_{\mathbb{Z}_{2}}{E_{2}})_{+}\,.$
Consider the composition
(5) $D_{2}(X_{+})@>c>>D_{2}(X)@>\hat{\mathbf{h}}_{2}>>Q(D_{2}(X))@>\iota_{X}>>Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}\,$,
where $\hat{\mathbf{h}}_{2}$ is the inclusion map.
If $((x,y),t)\in X^{\times 2}\times E_{2}$ is a point, we write $(x,y),t]\in X^{\times 2}\times_{\mathbb{Z}_{2}}E_{2}$ for the projection of $((x,y),t)$ to its $\mathbb{Z}_{2}$-orbit. Then (5) is given by
(6) $(x,y),t]\mapsto x\wedge y+y\wedge x\,$,
where again we write $x,y\colon S^{0}\to X$ for the based maps defined by the points $x,y\in X$.
Similarly, using definition of $H$, we see by direct calculation that the composition
(7) $D_{2}(X_{+})@>c>>D_{2}(X)@>\hat{\mathbf{H}}_{2}>>Q(D_{2}(X))@>\iota_{X}>>Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$
(in which $\hat{\mathbf{H}}_{2}$ is adjoint to ${\mathbf{H}}_{2}$) is given by
(8) $(x,y),t]\mapsto(x+y)\wedge(x+y)-((x\wedge x)+y\wedge y))\,.$
The map $D(X_{+})\to Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$ defined by $(x,y),t]\mapsto(x+y)\wedge(x+y)$ is homotopic to the sum of three terms
$(x,y),t]\mapsto x\wedge y+(x\wedge y+y\wedge x)+y\wedge y\,$,
where each of the terms defines a map $D(X_{+})\to Q_{\mathbb{Z}_{2}}(X\wedge X)^{\mathbb{Z}_{2}}$ in its own right. Therefore, (8) is homotopic to the signed sum of five maps
$(x,y),t]\mapsto x\wedge x+(x\wedge y+y\wedge x)+y\wedge y-x\wedge x-y\wedge y\,$,\\
and by homotopy commutativity, the latter is homotopic to the map (6). We infer that the (5) and (7) are homotopic.
As $\iota_{X}$ is a homotopy retract it follows that $\hat{\mathbf{H}}_{2}\circ c$ is homotopic to $\hat{\mathbf{h}}_{2}\circ c$, i.e., on homotopy classes $c^{\ast}(\hat{\mathbf{H}}_{2})=c^{\ast}(\hat{\mathbf{h}}_{2})$. Since $c^{\ast}$ is injective, we conclude that ${\mathbf{H}}_{2}$ is homotopic to ${\mathbf{h}}_{2}$. ∎
7. Proof of Addendum D
Let $T$ be the Quillen model category of compactly generated weak Hausdorff spaces [15]. A weak equivalence of $T$ is a weak homotopy equivalence. A fibration is a Serre fibration. A cofibration is a map which possesses the left lifting property with respect to the acyclic fibrations.
Let $T_{\ast}(\pi)$ be the category based left $\pi$-spaces and $\pi$-equivariant maps; this category has a zero object, i.e., the space $\ast$ consisting of a single point. Declare a map to be a weak equivalence (fibration) if and only if it is a weak equivalence (resp. fibration) upon application of the forgetful functor $T_{\ast}(\pi)\to T$. A map is a cofibration if it possesses the left lifiting property with respect to the acyclic fibrations. Then with respect to these choices, $T_{\ast}(\pi)$ is a model category by application of [7, thm. 11.6.1]. Note that the cofibrant objects of $T_{\ast}(\pi)$ are retracts of objects built up from a point by attaching free cells, where a free cell of dimension $j$ is $D^{j}\times\pi$. In particular, a free $\pi$-CW complex (in the base sense) is cofibrant. Note that every object is fibrant.
For cofibrant objects $X,Y\in T_{\ast}(\pi)$, let
$X,Y]_{\pi}=\hom_{\text{ho}T_{\ast}(\pi)}(X,Y)\,.$
Note that $Q_{\mathbb{Z}_{2}}(Y)$ has the structure of an object of $T_{\ast}(\pi\times\mathbb{Z}_{2})$. Suppose in addition $X$ has the structure of a based $(\pi\times\mathbb{Z}_{2})$-CW complex which is $\pi$-free after forgetting the $\mathbb{Z}_{2}$-action. In this case, we set
$\{X,Y\}_{\pi\times\underline{\mathbb{Z}_{2}}}:=\pi_{0}(\operatorname{map}_{T_{\ast}(\pi)}(X,Q_{\mathbb{Z}_{2}}(Y))^{\mathbb{Z}_{2}})\,$,
where $\operatorname{map}_{T_{\ast}(\pi)}(A,B)$ denotes the mapping space of based equivariant maps $A\to B$.
Proof of Addendum D.
With respect to the above conventions, the equivariant stable Hopf invariant is defined in the same way as when $\pi$ is trivial: For a cofibrant $\pi$-space $Y$, the tom Dieck fiber sequence
$\Sigma^{\infty}Y_{h\mathbb{Z}_{2}}\to(\Sigma^{\infty}_{\mathbb{Z}_{2}}Y)^{\mathbb{Z}_{2}}\to\Sigma^{\infty}Y^{\mathbb{Z}_{2}}$
is a fiber sequence of naive spectra with $\pi$-action which splits equivariantly making use of the inclusion $Y^{C^{2}}\to Y$ If we set $Y=B\wedge B$ where $B\in T_{\ast}(\pi)$ is a cofibrant object, we see then the construction of §4 yields
$H^{\pi}\colon\{A,B\}_{\pi}\to\{A,D_{2}(B)\}_{\pi}\,.$
The proof that properties (i)-(iv) hold is exactly the same as the proof of Theorem A above, where the maps and homotopy classes are to be interpreted equivariantly.
It remains to show that in the metastable range, a homotopy class $f\in\{A,B\}_{\pi}$ destabilizes if and only if $H^{\pi}(f)=0$. We will provide a crude sketch of the argument and leave the details to the reader. The Snaith splitting is natural, [^1] More precisely, the proofs that appear in [4] and [6] are natural. so the weak homotopy equivalence
$\Sigma^{\infty}Q(B)@>\simeq>>\bigvee_{n\geq 1}\Sigma^{\infty}D_{n}(B)$
is also $\pi$-equivariant. It follows that a $\pi$-equivariant version $h^{\pi}$ of the second Segal-Snaith Hopf invariant $h$ is defined. Moreover, the sequence
$B\to Q(B)@>h^{\pi}>>Q(D_{2}(B))$
is an equivariant fiber sequence in the metastable range by the Blakers-Massey excision theorem, where the map $h^{\pi}$ is induced by projection onto the second summand of the Snaith splitting. Then the equivariant stable Hopf invariant $h^{\pi}$ is the the effect of applying $A,{-}]_{\pi}$ to the map of the same name. We infer that $h^{\pi}$ yields the total obstruction to destabilization in the metastable range. Lastly, the proof we gave that $H=h$ is valid in the equivariant setting, and we conclude that $H^{\pi}=h^{\pi}$. ∎